Find the equation of the lines that goes through the given points (2,-4) and (-3,-9). Write the equation in both standard form and slope-intercept form.
Thanks.

Answer by Alan3354(69443)

Find the equation of the lines that goes through the given points (2,-4) and (-3,-9). Write the equation in both standard form and slope-intercept form.
----------------
The hard way:
find the slope
use the slope and either point to find the y-intercept
--> y = mx + b
==================
Use determinants:
----
|+x +y 1|
|+2 -4 1| = 0
|-3 -9 1|
Note: +x and +y are for alignment
---------
x*(-4+9) - y*(2+3) + 1*(-18-12) = 0
5x - 5y = 30
x - y = 6
===============
y = x - 6